A beer-bottling company ordered 2100 more 12-ounce bottles than 1-quart bottles. The cost of the 12-ounce bottles was $.08 each, and the cost of the 1-quart bottles was $.11 each. The total cost was $310.50 not including shipping. Find the number of each type of bottle ordered.
Answer by mananth(16946)

12 ounce bottles -------x
1 quart bottle-----------y
x-y=2100.............1
0.08x+0.11y=310.50
multiply by 100
8x+11y=31050.........2
1 x -1 y = 2100 .............1
8 x + 11 y = 31050 .............2
Eliminate y
multiply (1)by 11
Multiply (2) by 1
11 x -11 y = 23100
8 x 11 y = 31050
Add the two equations
19 x = 54150
/ 19
x = 2850
plug value of x in (1)
1 x + -1 y = 2100
2850 + -1 y = 2100
-1 y = 2100 -2850
-1 y = -750
y = 750